h2HW roblem 2.24 Part A Review How fast is the watermelon going when it passes Superman? Express your answer with the appropriate units As a science project, you drop a watermelon off the top of the Empire State Building, 320 m above the sidewalk. It so happens that Superman flies by at the instant you release the watermelon. Superman is headed straight down with a speed of 34.0 m/s 78 Submit Previous Answers Request Answer X Incorrect: Try Again; 4 attempts remaining Provide Feedback Type here to search 2 3
Answer #1

for water melon

initial velocity voy = 0


acceleration ay = -g


initial position yo = 320 m

at time t and position y

let the water melon passes the superman

y - yo = vy*t + (1/2)*ay*t^2

y - 320 = 0 - (1/2)*9.8*t^2 ............(1)


for super man


speed of v1 = -34 m/s


acceleration a1 = 0

initial position yo = 0

y - yo = v1*t + (1/2)*a1*t^2


y - 320 = -34*t...........(2)


from 1 and 2


-34*t = -(1/2)*9.8*t^2

time t = 6.94 m/s


at time t = 6.94 for water melon


vy = voy + ay*t

vy = 0 - 9.8*6.94 = -68.012 m/s

the speed of water melon = 68.012 m/s <<<<--------------answer